1) you observe a population with 4000 zygotes with the genotype AA,800 zygotes with the genotype Aa, and 50 zygotes with the genotype aa.

a) what are the frequencies of the three genotypes and the A and a alleles in the zygotes?

b) you come back later to this population and determined there are the following number of juveniles:3000 AA, 700 Aa, and 12 aa genotype. calculate the relative fitness w, and the selection coefficient for each of he 3 genotypes.

c) calculate normalized genotype frequencies after selection.

d) calculate allele frequencies after selection.

e) what kind of selection is occuring ad what will be the equilibrium allele frequencies?
Answer #1

1)

Based on the given data,

Zygotes have two types of alleles “A” and “a.” Combination of these two alleles would yield three types of genotypes: “AA,” “Aa” and “aa.” The given data is:

The Genotypic frequencies are:

  • AA = 4000 / 4850 = 82.47
  • Aa = 800 / 4850 = 16.49
  • aa = 50 / 4850 = 1.0

The analysed population has two alleles (A and a) at a locus, each zygote will have combination of these two alleles. Hence the whole population has 9700 (4850×2) alleles at a locus. To estimate frequency of alleles sum the number of A and a alleles, then divide by the total number of alleles.

Hence, the frequency of allele “A” is:

f(A) = [(2 x 4000) +800]/ 9700 = 0.9

f(a) = [(2 x50) +800]/ 9700 = 0.1

Thus, the allelic frequency of “A” is 0.9 and “a” is 0.1. (A + a = 1, hence it is a not polymorphic).